OpenStudy (anonymous):

I was doing this problem and the answer they give I think is wrong could some one run the problem and see what they get. 19/8 X 5/3

OpenStudy (anonymous):

95/24

OpenStudy (anonymous):

Yup, 95/24

OpenStudy (anonymous):

\[\Large \frac{{19}}{8} \times \frac{5}{3} = \frac{{19 \times 5}}{{8 \times 3}} = \frac{{95}}{{24}}\]

OpenStudy (anonymous):

yes but 95/24 =9 4/24 reduce 9 1/6 is that right?

OpenStudy (anonymous):

No. It can reduce to \(\Large 3\frac{{23}}{{24}}\) but I try to avoid mixed fractions at all costs.

OpenStudy (anonymous):

i get it now it was in my math thank you .